The USDA collects data on commodity prices and consumption. Use their data to explain per-capita chicken consumption as a function of the price of chicken. (Note that economic theory suggests that an increase in a commodity price is generally associated with a decrease in demand.) -per capita chicken consumption PCHICK consumer price index for chicken We estimated a linear regression model using annual data for 1950-2001 (N 52) Dependent variable: coefficient std. error const PCHICK -0.1 0.1 R-squared: 0.854880 a. Write the equation for the least-squares regression line: b. The null hypothesis is: (e)0 c. Calculate the t-ratio d. The critical value, evaluated at α = 0.05 , for the t distribution (with N-2 df) is e. Can we reject the null hypothesis? V (pick one) yes no
Answer #1

a)

least square regression line: Yhat=7+(-0.1)*x

b)

null hypothesis: Ho:\beta >=0

c)

t ratio =coefficient/std error =-0.1/0.1 =-1

d)critical value =-1.676

e)

No